Part A What is the resistance of a computer monitor that draws an rms current of...

Part A

What is the resistance of a computer monitor that draws an rms current of 0.437 A when connected to a 120-V outlet?

r=ohm?

Part B

What is the peak current of a computer monitor?

i0=?a

Part C

What is the power level of a computer monitor?

p=?w
Answer #1

a)

i = V/R = 120 / 0.437 = 274.6 ohm

b)

io = i sqrt (2) = 0.437* sqrt (2) = 0.618 A

c)

P = i V = 120 * 0.437 = 52.44 W

=======

Comment in case any doubt, will reply for sure.. Goodluck